# CoARA will not save science from the tyranny of administrative evaluation

### v2

@@ 1 Introduction
The agreement of the Coalition for Advancing Research Assessment (CoARA) is one of the key documents of the contemporary discussion on the role of research evaluation in the organization of science [Coalition for Advancing Research Assessment (2022), CoARA]. It comes after the so-called Leiden Manifesto on Research Assessment [Hicks et al., 2015], after the advocacy for the responsible use of metrics in evaluation [Wilsdon et al., 2015], and more than ten years after DORA, the San Francisco Declaration on Research Assessment [DORA, 2012]. All these documents can be considered as attempts by the “science reform movement” to respond to the long-lasting crisis of contemporary science [Peterson and Panofsky, 2020].
All these documents are mainly focused on research evaluation technologies, i.e. on the techniques used for realizing evaluation. CoARA promotes the widespread use of peer-review technology supported by the responsible use of metrics, as defined in the Leiden manifesto [Hicks et al., 2015, Wilsdon et al., 2015]. The current discussion against CoARA proposals is animated by the harsh criticism of scientometric advocates [Abramo, 2024] speaking of “scientometric denialism” [Torres-Salinas et al., 2023] or ridiculing the idea of narrative assessment [Torres-Salinas et al., 2024].
These documents and the related discussion relegate crucial issues to an invisible background such as the social desirability of evaluation, the costs and benefits associated with it, the institutional context within which it develops, and the fairness of the procedures that characterize it. They did not consider foundational questions such as the relations between research evaluation, democracy, academic freedom, pluralism, and the soundness of science (see, among others, [Gillies, 2008, Collini, 2012, De Nicolao, 2022, Caso, 2022, Pievatolo, 2024]
@@ 2 Administrative evaluation of research
“Administrative evaluation of research” encompasses massive research evaluation exercises such as the British Research Excellence Framework (REF), the Italian evaluation of research quality (VQR) or the Australian Excellence in Research (ERA); massive evaluations conducted at the individual level as in the case of the Spanish “sexenio” [Marini, 2018] or the qualification for professorship (ASN) in Italy.
Indicators and information obtained from administrative evaluation of research are usually considered the basis of evidence-based science and university policy and governance [Whitley, 2007].They are used for decisions on recruitment, promotions and awards for individual researchers, for the performance-based research funding (PBRF) systems [Zacharewicz et al., 2018], and for project funding.
@@
The use of “administrative evaluation of research” instead of the generic expression “research evaluation” allows to carefully distinguish it from the evaluation that scientists continuously carried out during their work. When a scientist reads a paper or a book and decides to recommend it for publication, or to use it in their own research, they implicitly judge it as worthy of use because of its soundness, originality, solidity, and other such qualities. When a scientific controversy arises, scientists evaluate and discuss the reasons of the opponents; “Scientific debate can continue indefinitely in search of a conclusion that can no longer be refuted" [Pievatolo, 2024, p. 6]. In contrast, in administrative evaluation of research, “assessment authorities” give evaluators decision-making power that allows them to decide when the debate should end and to impose their decisions “even on those who disagree”[Pievatolo, 2024, p. 6].
@@
In recent decades, administrative evaluation has become a “taken-for-granted aspect of public administration and organizational procedures for management and development”[Dahler-Larsen, 2011, p. 2]. administrative evaluation of research has also begun to be considered an inevitable aspect of research organization. This is the result of the widespread adoption of a consequentialist or utilitarian approach to administrative evaluation. In a nutshell, according to this vision, science has goals to achieve. These goals are defined by political authorities, supranational policies, or experts, perhaps after consultation with stakeholders. The scientific system is essentially designed as a system for producing science. The only relevant feature of the scientific system is the efficiency of production; hence, socially desirable scientific institutions are efficient institutions. The allocation of resources to these institutions should aim to maximize the expected outcomes. Research evaluation is functional in enhancing the efficiency of research institutions; it allows political authorities or governing bodies to gather information to steer research institutions at distance. In particular, it allows one to properly allocate resources to the institutions that will be able to make the best use of them. In turn, this mechanism generates a continuous push toward improving research, enhancing the efficiency of the research system, and the growth of the scientific wealth of nations [Whitley, 2007, Dahler-Larsen, 2011, Davis et al., 2012, Hammarfelt and Hallonsten, 2022]. This approach is fully defined inside a model that argues that the quality and efficiency of science institutions are improved by introducing management techniques and practices that mimic the functioning of the market. According to Bleiklie [2018], this model was the result of the recipes of the new public management, and it can be labelled as “neoliberal” (see also [Lave et al., 2010, Mirowski, 2011]).
@@ 3 CoARA and the promotion of evaluative technocracy
As anticipated, the domain of CoARA is administrative evaluation of research. In fact, “administrative evaluation of research” coincides with the CoARA definition of “research assessment” encompassing “the assessment of research performing organizations and research units, by assessment authorities, research funding and performing organizations; […] The assessment of research projects by assessment authorities, research funding and performing organizations, and prize awarding organizations; […] The assessment of individual researchers and research teams by research funding and performing organizations and prize awarding organizations” [Coalition for Advancing Research Assessment (2022), CoARA].
CoARA considers administrative evaluation of research as a taken-for-granted aspect of the contemporary organization of science, by avoiding the discussion about its legitimacy and neutrality with respect to the goal of science. More precisely, CoARA explicitly assumes that administrative evaluation of research is desirable for science and society: “the assessment of research, researchers and research organizations recognizes the diverse outputs, practices and activities that maximize the quality and impact of research” [Coalition for Advancing Research Assessment (2022), CoARA, p. 1]. CoARA contributes to the administrative evaluation of research by simply adding that the adoption of the most robust evaluation technique is essential for efficiency, i.e. to “maximize the quality and impact of research”. According to CoARA, “peer review is the most robust method known for assessing quality and has the advantage that it is in the hands of the research community” [Coalition for Advancing Research Assessment (2022), CoARA, p. 5].
@@
CoARA agreement is exclusively addressed to “research funding organizations, research performing organizations, national/regional assessment authorities and agencies” [Coalition for Advancing Research Assessment (2022), CoARA, p.2 ]. COARA does not have a direct approach to researchers in the form of a request for the signing of deontological commitments. They are primarily the subjects to be assessed. Their contribution to reform consists of supporting the dissemination and adoption of COARA principles within their institutions. In this respect, CoARA adopts an approach very different from that of DORA. DORA is aimed at a broader target audience, including researchers who, by signing the commitments, agree to act in a deontological manner, adopting specific behaviors in research and when acting as evaluators [DORA, 2012, see recommendations 15-18].
@@
Indeed, CoARA’s four core commitments concern the technologies to be adopted or avoided for administrative evaluation. The first consists in the recognition in administrative evaluation of the diverse outputs, practices, and activities that reflect the variety of contributions to science. The second requires that administrative evaluation be based “on qualitative evaluation for which peer review is central, supported by responsible use of quantitative indicators”. This second requirement incorporates the specific actions suggested in the Leiden manifesto for use of metrics [Hicks et al., 2015], explicitly recalled in Annex 4 of the agreement. The third commitment supports the abandonment of “inappropriate uses of journal- and publication-based metrics, such as journal impact factors and h-index”. It explicitly recalls, again in Annex 4, the specific actions described in the DORA declaration [DORA, 2012]. The fourth commits the signatories of the agreement to “avoid the use of rankings of research organizations in research assessment” [Coalition for Advancing Research Assessment (2022), CoARA, pp. 4-6].
@@
The other six supporting commitments “include three commitments to enable the move towards new research assessment criteria, tools and processes, and three commitments to facilitate mutual learning, communicate progress and ensure that new approaches are evidence informed” [Coalition for Advancing Research Assessment (2022), CoARA, p. 7].
@@
In general, the ten commitments adopt an ‘internalist’ or technological’ approach to administrative evaluation of research and its reform. In a nutshell, the reform of the administrative evaluation of research proposed by CoARA consists in a change of technology for evaluation. It is possible to express this point with the metaphor of the thermometer, At least since Moed et al. [1985], the thermometer represents bibliometrics, i.e. the (best) technology used for quantitative research evaluation. CoARA proposes to replace this technology, the bibliometric thermometer, with another one, the peer review, supported by the responsible use of quantitative indicators.
@@
CoARA lacks of a precise definition of “peer review” and a discussion of the many types of “peer review” adopted in contemporary science. It appears that CoARA promotes the adoption of qualitative evaluations by the peer community, regardless of the specific ways in which they are performed. The replacement of metrics with peer review assumes, in fact, that when researchers act as evaluators, they are, as scientists, naturally able to recognize research quality and impact. CoARA admits that evaluators can be consciously or unconsciously biased on gender, ethnic origin, and other, and encourages training evaluators to recognize and mitigate biases [Coalition for Advancing Research Assessment (2022), CoARA, p. 22]. It is implicit in CoARA that “unbiased” evaluators are committed to the good of science, independently of the institutional and administrative framework in which the assessment is realized. This appears hard to defend, as it asks scientists not to be self-interested. A more realistic way of representing the behavior of scientists engaged in administrative evaluation consists in admitting that they may act as “rational referees, who might not have any incentive to see high-quality work other than their own” [Thurner and Hanel, 2011] or of their group or institution. This is all the more true since the issues involved are mostly about money (e.g., hiring, promotions, funding). Agent-based simulations on peer-review in journals show that the presence of a minority of rational or selfish referees is sufficient to drastically lower the quality of the scientific standard [Thurner and Hanel, 2011, Cabotà et al., 2013]. CoARA appears to hide or ignore such problems.
This replacement of technologies also assumes that there are no differences between the evaluation of research carried out within an administrative process, such as the hiring of a researcher or the national research evaluation, and the common evaluative practices adopted by scholarly communities to recognize scientific claims as provisionally true [Kitcher, 2011, Pievatolo, 2024]. More precisely, CoARA operates in the slippery landscape as defined by the argument that peer review “is in the hands of the research community” [Coalition for Advancing Research Assessment (2022), CoARA, p. 5]: the judgments of scientists involved in the administrative evaluation of research are of the same nature or type as those produced in the peer review processes of academic journals [Bornmann, 2011], and as those discussed by scientists in public debates and controversies about science. In fact, all these evaluative practices use different information, possibly adopt different standards, and happen in different contexts. COARA hides or ignores such differences.

### v3

@@ 1 Introduction
The agreement of the Coalition for Advancing Research Assessment (CoARA) is one of the key documents of the contemporary discussion on the role of research evaluation in the organization of science [CoARA, 2022]. It comes after the so-called Leiden Manifesto on Research Assessment [Hicks et al., 2015], after the advocacy for the responsible use of metrics in evaluation [Wilsdon et al., 2015], and more than ten years after DORA, the San Francisco Declaration on Research Assessment [DORA, 2012]. All these documents, with others (for a syntesis see [Rushforth, 2025]), can be considered as attempts by the “science reform movement” to respond to the long-lasting crisis of contemporary science [Peterson and Panofsky, 2020].
All these documents are mainly focused on research evaluation technologies, i.e. on the techniques used for realizing evaluation. CoARA promotes the widespread use of peer-review technology supported by the responsible use of metrics, as defined in the Leiden manifesto [Hicks et al., 2015, Wilsdon et al., 2015]. The use of peer review is instrumental in shifting the focus from traditional metrics, such as publication counts and citation indices, to a broader appreciation of diverse academic contributions, including teaching, leadership, openness, and societal impact. The debate between reformers and advocates of metrics is ongoing. In the Netherlands, for example, proponents of the Recognition and Rewards Program, considered the main ongoing experiment in assessment reform, face opposition from professors who express concern about the potential risks, especially for young scholars, arising from ’the abandonment of international evaluation criteria’ [Rushforth, 2025]. The tone of the debate is much more heated with regard to CoARA directly. The primary criticisms come from advocates of scientometrics [Abramo, 2024], who even accuse CoARA of ’scientometric denialism’ [Torres-Salinas et al., 2023]. Critics did not spare even the narrative curriculum, one of the main instruments proposed by CoARA. It consists of a written narrative that contextualizes a researcher’s contributions, explaining the quality, relevance, and impact of their work in a broader sense—including teaching, leadership, open science, and societal impact. Torres-Salinas et al. [2024] ridiculized the narrative curriculum by launching the idea of “narrative bibliometrics”, and illustrating how to write a narrative bibliometric curriculum, by using word to contextualize metrics .
Reform documents and the related discussions relegate crucial issues to an invisible background such as the social desirability of evaluation, the costs and benefits associated with it, the institutional context within which it develops, and the fairness of the procedures that characterize it. In particular, they did not consider foundational questions such as the relations between research evaluation, democracy, academic freedom, pluralism, and the soundness of science (see, among others, [Gillies, 2008, Collini, 2012, De Nicolao, 2022, Caso, 2022, Pievatolo, 2024a]
@@ 2 Administrative evaluation of research
Indicators and information obtained from administrative research evaluation are typically regarded as the fundamental components of accountability for the use of taxpayer funds in research. They also serve for evidence-based science policy and university governance [Whitley, 2007]: they are instrumental in decisions regarding recruitment, promotions, and awards for individual researchers, as well as for project funding and performance-based research funding (PBRF) systems [Zacharewicz et al., 2018]. Molas-Gallart [2012] proposed to distinguish three main uses of the administrative research evaluation: distributive, aimed at distributing resources according to the results of evaluation; improvement, aimed at individuating the best practice; and controlling, aimed at scritinizing the use of public resurces. As a matter of fact “administrative evaluation of research” encompasses massive research evaluation exercises such as the British Research Excellence Framework (REF), the Italian evaluation of research quality (VQR) or the Australian Excellence in Research (ERA); the performance based funding systems in the European nordic countries [Aagaard et al., 2015], massive evaluations conducted at the individual level as in the case of the Spanish “sexenio” [Marini, 2018] or the qualification for professorship (ASN) in Italy; evaluation of projects for funding as in the European Research Council competition for ERC grant. Indeed, research governance structures vary across countries, leading to diverse applications and uses of administrative research evaluation.
@@
The espression “administrative evaluation of research” instead of “research evaluation” allows to carefully distinguish it from the evaluation that scientists continuously carried out during their work. When a scientist reads a paper or a book and decides to recommend it for publication, or to use it in their own research, they implicitly evaluate it as worthy of use because of its soundness, originality, solidity, and other such qualities. When a scientific controversy arises, scientists evaluate and discuss the reasons of the opponents; “Scientific debate can continue indefinitely in search of a conclusion that can no longer be refuted" [Pievatolo, 2024a, p. 6]. In contrast, in administrative evaluation of research, “assessment authorities” give evaluators decision-making power that allows them to decide when the debate should end and to impose their decisions “even on those who disagree”[Pievatolo, 2024a, p. 6].
@@
In recent decades, administrative evaluation has become a “taken-for-granted aspect of public administration and organizational procedures for management and development”[Dahler-Larsen, 2011, p. 2]. Hence, the administrative evaluation of research has also begun to be considered an inevitable aspect of research organization, despite the enormous variety of the ways in which it is used in different countries. This is the result of the widespread adoption of a consequentialist or utilitarian approach to administrative evaluation. In a nutshell, according to this vision, science has goals to achieve. These goals are defined by political authorities, supranational policies, or experts, perhaps after consultation with stakeholders. The scientific system is essentially designed as a system for producing science. The only relevant feature of the scientific system is the efficiency of production; hence, socially desirable scientific institutions are efficient institutions. The allocation of resources to these institutions should aim to maximize the expected outcomes. Administrative research evaluation is functional in enhancing the efficiency of research institutions; it allows political authorities or governing bodies to gather information to steer research institutions at distance. In particular, it allows one to properly allocate resources to the institutions that will be able to make the best use of them. In turn, this mechanism generates a continuous push toward improving research, enhancing the efficiency of the research system, and the growth of the scientific wealth of nations [Whitley, 2007, Dahler-Larsen, 2011, Davis et al., 2012, Hammarfelt and Hallonsten, 2022]. This approach is fully defined inside a model that argues that the quality and efficiency of science institutions are improved by introducing management techniques and practices that mimic the functioning of the market. According to Bleiklie [2018], this model was the result of the recipes of the new public management, and it can be labelled as “neoliberal” (see also [Lave et al., 2010, Mirowski, 2011, Kulczycki, 2023]).
@@ 3 CoARA and the promotion of evaluative technocracy
In March 2021 the European Commission (EC), Directorate for Research and Innovation, Unit A.4 Open Science, started a consultation with “European and international stakeholders on how to facilitate and speed up reform so that the quality, performance and impact of research and researchers are assessed on the basis of more appropriate criteria and processes” [European Commission et al., 2021, p. 3]. The EC considered that the current research assessment system "uses inappropriate and narrow methods to assess the quality, performance and impact of research and researchers. Notably, the quantity of publications in journals with high Journal Impact Factor and citations are currently the dominant proxies for quality, performance and impact". Consultation results were synthesized in a scoping report published in November 2021. It proposed the writing “of a European agreement that would be signed by individual research funding organisations, research performing organisations and national/regional assessment authorities and agencies, as well as by their associations, all willing to reform the current research assessment system” [European Commission et al., 2021, p. 3]. The text of the Agreement on Reforming Research Assessment (ARRA) was written from February to July 2022. The writing of the ARRA was conducted by a drafting team of 12 people from European Commission and Science Europe, a non-profit international organization gathering forty national research funding organizations. A group of 19 “contributors” participated in the writing through comments and suggestions [Arentoft et al., 2022]. The draft of ARRA was also discussed in three assemblies involving organizations that had participated in the consultations for the scoping report. Finally, the member states of the European Union were also consulted (for a more detailed reconstruction of the whole process, see Di Donato [2024]). On 28 September 2022 the agreement was officially opened for signature and CoARA, the Coalition for the Advancement of Research Assessment, was launched. CoARA is the “collective organization” in charge of putting the ARRA principles and commitments into practice. The modalities of governance, operations and funding of CoARA was drafted by “representatives from the European University Association, Science Europe and the European Commission” with the advice of an implementation group “made of representative” from ten organizations [CoARA, 2022]. On December 1st 2022 CoARA was officially launched during a “Constitutive assembly” that adopted the governance document. Actually, CoARA’s Secretariat is hosted by the European Science Foundation, a non-governmental, internationally-oriented, non-profit association established in France in 1974. It coordinates also the CoARA Boost Horizon Europe project, a project financed with 5 million Euros by the European Commission [CoARA, 2023b].
The European Commission presented ARRA as “a co-creation process involving more than 350 organizations from 40 countries” [Commission, 2022]. In a recent European Union report the whole process is summarized as “the [European] Commission facilitated a co-creation process to reach an agreement on reforming research assessment” [Oancea, 2024, p. 13]. Matthias Koenig, one of the current member of the steering board, even describes CoARA as “a science-driven, bottom-up and inclusive coalition” [CoARA, 2025].
However, despite the appearance of broad consultation and stakeholder involvement, the key orientations, procedures, and institutional architecture of CoARA were all designed and directed by the European Commission, with the support of selected “representative organizations”. Notably, even “representative organizations” and the stakeholders involved in the consultation were possibly chosen by the very institutions responsible for drafting the agreement. The rationale on the basis of which drafters, contributors, organisations and stakeholders are selected is not disclosed, nor is the information on from whom they receive mandates, funding or reimbursement of expenses, to whom they are accountable for decisions made [Pinto, 2023]. In sum, the text of ARRA, the governance structure, and operational modalities of CoARA were predefined by a restricted group of institutional actors “facilitated” by EC. Walter Rosenthal, the Vice president of the German Rector’s Conference, stated syntetically that CoARA “is basically the result of a top-down process” [Naujokaitytė, 2023].
As suggested in the scoping report, CoARA agreement is exclusively addressed to “research funding organizations, research performing organizations, national/regional assessment authorities and agencies” [CoARA, 2022, p.2 ]. Once an organization has signed the agreement or joined CoARA as a member, signatory institutions and members of CoARA are granted operational flexibility in how they implement the commitments undertaken: organization have to present an action plan within a year of signing the agreement, and they “have full freedom in the develoment of their Action Plan” [CoARA, 2023a]. At April 2025, 191 action plans are available in Zenodo repository (https://coara.eu/agreement/action-plan/).
@@
The responsibility for translating CoARA’s principles into concrete practices rests entirely with the member institutions and organization. Indeed, CoARA does not directly engage with individual researchers: for example, it does not require them to sign any form of ethical or deontological commitment. Researchers’s contribution to reform consists only of supporting the dissemination and promoting the adoption of COARA principles within their institutions. In this respect, CoARA differs from DORA that is aimed at a broader target audience, including researchers who, by signing its commitments, agree to act in a deontological manner, adopting specific behaviors in research and when acting as evaluators [DORA, 2012, see recommendations 15-18].
The domain of CoARA is solely “research assessment”, the definition of which coincides with that of administrative evaluation of research. According to CoARA, “research assessment” encompasses “the assessment of research performing organizations and research units, by assessment authorities, research funding and performing organizations; […] The assessment of research projects by assessment authorities, research funding and performing organizations, and prize awarding organizations; […] The assessment of individual researchers and research teams by research funding and performing organizations and prize awarding organizations” [CoARA, 2022].
CoARA considers administrative evaluation of research as a taken-for-granted aspect of the contemporary organization of science, by avoiding the discussion about its legitimacy and neutrality with respect to the goal of science. More precisely, CoARA explicitly assumes that administrative evaluation of research is desirable for science and society: “the assessment of research, researchers and research organizations recognizes the diverse outputs, practices and activities that maximize the quality and impact of research” [CoARA, 2022, p. 1]. CoARA contributes to the administrative evaluation of research by simply adding that the adoption of evaluations techniques coherent with the principles of the agreement is essential for efficiency, i.e. to “maximize the quality and impact of research”. According to CoARA, “peer review is the most robust method known for assessing quality and has the advantage that it is in the hands of the research community” [CoARA, 2022, p. 5].
@@
Indeed, CoARA’s four core commitments concern the technologies to be adopted or avoided for administrative evaluation. The first consists in the recognition in administrative evaluation of the diverse outputs, practices, and activities that reflect the variety of contributions to science. The second requires that administrative evaluation be based “on qualitative evaluation for which peer review is central, supported by responsible use of quantitative indicators”. The third commitment supports the abandonment of “inappropriate uses of journal- and publication-based metrics, in particular inappropriate uses of Journal Impact Factor (JIF) and h-index”. The fourth commits the signatories of the agreement to “avoid the use of rankings of research organizations in research assessment” [CoARA, 2022, pp. 4-6].
@@
The other six supporting commitments “include three commitments to enable the move towards new research assessment criteria, tools and processes, and three commitments to facilitate mutual learning, communicate progress and ensure that new approaches are evidence informed” [CoARA, 2022, p. 7].
It is worthwhile to note that CoARA, particularly its second and third commitments, does not incorporate specific actions such as those proposed in DORA [DORA, 2012] or in the Leiden Manifesto for research metrics [Hicks et al., 2015]. Both documents are explicitly referenced only in Annex 4 of the Agreement, but the annexes “do not form an integral part of the Agreement” and are “provided to support its implementation” [CoARA, 2022, p. 12]. Moreover, as noted above, the Agreement grants signatories full discretion in selecting the tools and approaches to be adopted in their action plans. As a result, these plans may diverge significantly from the specific principles laid out in DORA and the Leiden Manifesto (see for instance [Pievatolo, 2024b]). This discretion also extends to open science and open access, which are neither mandated nor operationalized in a uniform or binding way. The definition of “openness” provided in the Agreement is rather narrow: “openness corresponds to early knowledge and data sharing, as well as open collaboration including societal engagement where appropriate” [CoARA, 2022, p. 3]. Signatories are only committed to rewarding “research behaviour underpinning open science practices such as early knowledge and data sharing as well as open collaboration” [CoARA, 2022, p. 4]. In contrast, the broader and more robust principles of the UNESCO recommendation on open science [UNESCO, 2021] are relegated to Annex 4, while open access is mentioned only in Annex 1. Given this framework, it is perhaps unsurprising that action plans might be developed with little or no attention to open science or open access [Galimberti, 2024].
@@
In general, the ten commitments adopt an ‘internalist’ or technological’ approach to administrative evaluation of research and its reform. In a nutshell, the reform of the administrative evaluation of research proposed by CoARA consists in a change of technology for evaluation. It is possible to express this point with the metaphor of the thermometer. At least since Moed et al. [1985], bibliometrics is represented as a thermometer, i.e. the (best) technology used for quantitative research evaluation. CoARA proposes to replace this technology, the bibliometric thermometer, with another one, the peer review, supported by the responsible use of quantitative indicators.
@@
CoARA lacks of a precise definition of “peer review” and a discussion of the many types of “peer review” adopted in contemporary science. It appears that CoARA promotes the adoption of qualitative evaluations by the peer community, regardless of the specific ways in which they are performed. The replacement of metrics with peer review assumes, in fact, that when researchers act as evaluators, they are, as scientists, naturally able to recognize research quality and impact. CoARA admits that evaluators can be consciously or unconsciously biased on gender, ethnic origin, and other, and encourages training evaluators to recognize and mitigate biases [CoARA, 2022, p. 22]. It is implicit in CoARA that “unbiased” evaluators are committed to the good of science, independently of the institutional and administrative framework in which the assessment is realized. This appears hard to defend, as it asks scientists not to be self-interested. A more realistic way of representing the behavior of scientists engaged in administrative evaluation consists in admitting that they may act as “rational referees, who might not have any incentive to see high-quality work other than their own” [Thurner and Hanel, 2011] or of their group or institution. This is all the more true since the issues involved are mostly about money (e.g., hiring, promotions, funding). Agent-based simulations on peer-review in journals show that the presence of a minority of rational or selfish referees is sufficient to drastically lower the quality of the scientific standard [Thurner and Hanel, 2011, Cabotà et al., 2013]. CoARA appears to hide or ignore such problems.
This replacement of technologies also assumes that there are no differences between the evaluation of research carried out within an administrative process, such as the hiring of a researcher or the national research evaluation, and the common evaluative practices adopted by scholarly communities to recognize scientific claims as provisionally true [Kitcher, 2011, Pievatolo, 2024a]. More precisely, CoARA operates in the slippery landscape as defined by the argument that peer review “is in the hands of the research community” [CoARA, 2022, p. 5]: the judgments of scientists involved in the administrative evaluation of research are of the same nature or type as those produced in the peer review processes of academic journals [Bornmann, 2011], and as those discussed by scientists in public debates and controversies about science. In fact, all these evaluative practices use different information, possibly adopt different standards, and happen in different contexts. COARA hides or ignores such differences.